Dinitrogen tetroxide decomposes to nitrogen dioxide. Write the balanced chemical equation for this reaction. Start with pure N2O4 at 1.00 atm. Find the equilibrium partial pressures of N2O4 and NO2. The equilibrium constant for this reaction is K = 11.
Answer #1

                                                                        N2O4 (g) 2 NO2 (g)

Initial pressure                                         1.00                0

Change in pressure                                          - X                 2X

Equilibrium pressure                                    (1.00 - X)            2X

or, 11 = (2X)2/(1.00 - X)

or, 11 = 4X2/(1.00 - X)

or, 4X2 + 11X - 11 = 0

or,

or,

or,

or, X = 0.775 atm, - 3.53 atm

The acceptable value of X = 0.775 atm as partial pressure can not be negative.

Therefore, the equilibrium partial pressure of N2O4 = (1.00 - 0.775) = 0.225 atm

And the equilibrium partial pressure of NO2 = 2 x 0.775 = 1.55 atm